A 653 W Carnot engine operates between constant-temperature reservoirs at 133°C and 63.8°C. What is the rate at which energy is (a) taken in by the engine as heat and (b) exhausted by the engine as heat?
Answer #1

a)
P = 653 W, this is power coming out of engine

effeienecy = 1 - (TL/TH)
= 1 - ((63.8+273)/(133+273))
= 1 - (336.8)/(406)
= 0.17044

Pout/Pin = effeciency
653/Pin = 0.17044
Pin = 3831.2 W
Answer: 3831.2 W

b)
Rate of hear exhausted = Pin-Pout
= 3831.2 - 653
= 3178.2 W
Answer: 3178.2 W